OpenStudy (anonymous):

Graph the line of the equation 5x + 3y =-15 ow did you know to put -5 and - 3 ? did you change it because the answer equals -15?

jimthompson5910 (jim_thompson5910):

when x = 0, y is 5x + 3y = -15 5(0) + 3y = -15 0 + 3y = -15 3y = -15 y = -15/3 y = -5 So the y-intercept is (0, -5)

jimthompson5910 (jim_thompson5910):

To find the x-intercept, you plug in y = 0 and solve for x

OpenStudy (anonymous):

Thanks i'll try my best!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

would you use y=mx + b? in this situation?